A chemist needs to mix an 18% acid solution with a 45% acid solution to obtain a 12 liter mixture consisting of 36% acid. How many liters of each acid solutions must be used?
Answer #1
amount of 18% solution --- x L
amount of 45% solution --- 12-x L
.18x + .45(12-x) = .36(12)
times 100
18x + 45(12-x) = 36(12)

take it from here
answered by: Writeacher